OpenStudy (evanhelps):

Find the standard form of the equation of the parabola with a focus at (0, -9) and a directrix y = 9

hero (hero):

Insert points focus (0,-9) and directrix (x,9) in to the following formula: \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\) Upon doing so you get: \((x - 0)^2 + (y - (-9))^2 = (x - x)^2 + (y - 9)^2\) Which simplifies to: \(x^2 + (y + 9)^2 = (y - 9)^2\) And expands to: \(x^2 + y^2 + 18y + 81 = y^2 - 18y + 81\) \(y^2\) and \(81\) cancels on both sides leaving just: \(x^2 + 18y = -18y\) \(x^2 = -36y\) From here, you can isolate y to find the standard form of the equation of the parabola.

OpenStudy (evanhelps):

Oh so y=-1/36x^2

OpenStudy (evanhelps):

Thanks man I understand now